# Few-Shot Learning on Graphs: from Meta-learning to Pre-training and Prompting

### v3

@@ 2 Meta-learning Approaches
Prevailing episodic meta-learning methods are designed to learn a prior from a series of meta-training tasks, which is transferable to the downstream meta-testing tasks. The concept is illustrated in Fig. 2, using few-shot node classification on a graph as an example. In this scenario, meta-training tasks are constructed using base classes with ample labels (“base set”), whereas a meta-testing task involves novel classes with only few-shot labels. Each task comprises a support set and a query set, which function as training and testing data, respectively, within each task Finn et al. (2017). The support set is always labeled, whereas the query set is only labeled in meta-training but is unlabeled for inference and prediction in meta-testing. The prior is optimized on the meta-training tasks to enable rapid, lightweight adaptation using the support set that can minimize the loss on the query set. Hence, for the prior to be transferable to meta-testing tasks, such meta-learners assume that the meta-training and -testing tasks are sampled from an i.i.d. task distribution.
To tailor to few-shot learning on graphs, standard meta-learning methodologies have been enhanced in various ways. We further categorize the literature according to their graph-specific enhancements within the meta-learning framework, namely, structure-based enhancement and adaptation-based enhancement, as summarized in Table 1.
Method Enhancement Meta-learner Task
Node Edge Graph
GPN Ding et al. (2020) structure (node) Protonets $\checkmark$ $\times$ $\times$
FAAN Sheng et al. (2020) structure (node) Protonets $\times$ $\checkmark$ $\times$
HMNet Xiao et al. (2021) structure (edge) MN $\times$ $\checkmark$ $\times$
Wang et al. (2019b) structure (edge) MAML $\times$ $\checkmark$ $\times$
RALE Liu et al. (2021b) structure (path) MAML $\checkmark$ $\times$ $\times$
G-Meta Huang and Zitnik (2020) structure (subgraph) hybrid $\checkmark$ $\checkmark$ $\times$
GEN Baek et al. (2020) structure (subgraph) Protonets $\times$ $\checkmark$ $\times$
GFL Yao et al. (2020) adaptation (graph) MAML $\checkmark$ $\times$ $\times$
MI-GNN Wen et al. (2021) adaptation (graph) hybrid $\checkmark$ $\times$ $\times$
MetaTNE Snell et al. (2017) adaptation (task) Protonets $\checkmark$ $\times$ $\times$
AS-MAML Ma et al. (2020) adaptation (step size) MAML $\times$ $\times$ $\checkmark$
Table 1: Meta-learning approaches for few-shot learning on graphs.
2.1 Structure-based enhancement
@@
Node-level enhancement. The nodes of a graph encompass implicit structural information for enhancing the meta-learning model. Recent studies predominantly focus on differentiating node weights in a task to reflect their varying structural importance. Specifically, in Protonets-based methods such as GPN Ding et al. (2020) and FAAN Sheng et al. (2020), it is observed that each support node contributes variably to the prototype of each class in a task. Thus, they propose to evaluate the importance of nodes in the support set and adjust their contribution w.r.t. the query node when constructing the prototypes.
@@
Edge-level enhancement. Edges are fundamental to graphs as they represent explicit first-order structures. Moreover, a sequence of edges forms a path, which is capable of representing more complex high-order interactions among distant nodes. As a result, many meta-learning methods on graphs incorporate edge- or path-based enhancements. On one hand, edge-based enhancement leverages auxiliary information associated with edges, thereby extending beyond mere node-level information. For instance, to boost few-shot link prediction in knowledge graphs, HMNet Xiao et al. (2021) adopts matching networks (MN) as its meta-learning framework, performing matching for both entities and relations. In particular, for relation-level matching, it calculates a relation-aware score using a relation encoder to enrich the link prediction task. Another work Wang et al. (2019b) develops a generative model to generate new triplets, serving to augment the sparse links within a MAML-based framework. On the other hand, paths can capture long-range dependencies between distant nodes. RALE Liu et al. (2021b) proposes to exploit the paths between each query node and the support nodes for task-level dependencies, and between each target node and the hub nodes for graph-level dependencies.
@@
Subgraph-level enhancement. Subgraphs provide rich contextual information to its containing nodes. Hence, recent studies have utilized subgraphs to enhance the expressiveness of nodes. G-Meta Huang and Zitnik (2020) generates class prototypes based on the contextual subgraph of each support node, and similarly expands each query node to its subgraph, to enhance the robustness of prototype and query representations. Similarly, GEN Baek et al. (2020) proposes a graph extrapolation network to extrapolate knowledge through the neighbors of the support set, which essentially exploits the one-hop subgraphs of the support nodes.
2.2 Adaptation-based enhancement
@@
GFL Yao et al. (2020) recognizes the topological variances across different graphs, and thus customizes a global prior for each individual graph. More concretely, it generates class prototypes specifically tailored to each graph, utilizing a gate function based on graph-specific representation that is applied to the global prior. MI-GNN Wen et al. (2021) adopts a similar idea and employs a hypernetworks Perez et al. (2018) to modulate the global prior for each graph before undergoing a further task-wise adaptation within a MAML framework. Meanwhile, there is potential for enhancement in task-wise adaptation. In the context of multi-label few-shot classification, MetaTNE Lan et al. (2020) observes an increase in task diversity due to the multi-label setting. Thus, adopting a uniform node embedding across such diverse tasks may result in inadequate expressiveness, given that the same node could be associated with different labels in different tasks. To address this, MetaTNE conducts an additional adaptation for the node embeddings w.r.t. the query set in each task, thereby better adapting to the characteristics of the task. In addition to enhancing adaptation from the perspective of graph or task, AS-MAML Ma et al. (2020) improves adaptation from an optimization standpoint. It employs a reinforcement learning-based controller to determine the optimal step size for the adaptation process in MAML.
2.3 Discussion
@@
In summary, existing research for few-shot learning on graphs often enhances a standard meta-learner, either through structural augmentation or by refining the adaptation process. However, these approaches share a common drawback inherent to meta-learning approaches. First, they still require a large volume of labels for a “base set” during the meta-training phase, which can be challenging or costly to acquire. Second, they fail to leverage the vast amount of unlabeled data to learn a more comprehensive prior. Third, they are limited by the assumption in task distribution, and cannot handle different types of tasks downstream. This leads to a further challenge: How can we address a diverse range of few-shot learning tasks on graphs without an extensively annotated base set, while utilizing abundant unlabeled graph data? Recently, pre-training on graphs has emerged as a promising solution to this challenge.
3 Pre-training Approaches
@@
Pre-training for few-shot learning on graphs typically utilizes self-supervised methods on unlabeled graph data to train a graph encoder, which aims to capture task-agnostic intrinsic knowledge in graphs, in the pre-training stage. Typical knowledge include as local and global structures, as well as node features. The pre-trained graph encoder, serving as a form of prior knowledge, can be used to address a variety of downstream tasks through an adaptation stage. An overview of pre-training and adaptation process is depicted in Fig. 3.
Given the abundance of unlabeled graph data in many domains and the flexibility to deal with many types of downstream tasks, pre-training approaches have been popularized. In this section, we provide a taxonomy of existing work in this direction, covering both the pre-training stage and the downstream adaptation stage.
@@
Method Target Pos/Neg samples Augmentation
instance
GRACE Zhu et al. (2020) node node/node random
GCC Qiu et al. (2020) subgraph subgraph/subgraph random
GraphCL You et al. (2020) graph graph/graph random
DGI Veličković et al. (2018) graph node/graph random
InfoGraph Sun et al. (2019) graph node/graph random
MVGRL Hassani and Khasahmadi (2020) subgraph node/subgraph diffusion
JOAO You et al. (2021) graph graph/graph automated
Table 2: Summary of contrastive pre-training strategies.
3.1 Strategies for the pre-training stage
@@
Current self-supervised tasks (also known as pretext tasks) for graph pre-training predominantly fall into two broad categories: contrastive and generative strategies.
@@
Contrastive strategies. Contrastive pre-training functions by contrasting instances at various scales within a graph. For each target instance, contrastive methods sample positive and negative instances, aiming to bring the positive instances closer to the target in the latent space while distancing the negative instances. A summary of contrastive strategies is presented in Table 2.
@@
First, the scale of instances used for contrasting varies, ranging from individual nodes to broader subgraphs, or even the entire graph. One one hand, node-level approaches provide a more granular treatment focusing on the representations of each individual node. An example is GRACE Zhu et al. (2020), which generates two augmented views by randomly corrupting the original graph. Then, for each target node in one view, it treats the corresponding node from the other view as positive, while others as negative. On the other hand, graph-level approaches expand the scope to (sub)graphs, thereby capturing global structural patterns. GCC Qiu et al. (2020) conducts random walk in an $r$-hop ego network to obtain a target subgraph instance alongside its positive instances. Negative instances are acquired by performing random walk in an $r^{\prime}$-hop ego network of the same ego node, for some $r\neq r^{\prime}$. Similarly, GraphCL You et al. (2020) generates graph views with various data augmentations and treats graph views augmented from the same graph as positive instances, while those from different graphs as negative instances. Furthermore, there exist cross-scale approaches contrasting between node- and graph-level instances, such as DGI Veličković et al. (2018) and InfoGraph Sun et al. (2019). They both maximize the mutual information between the local node representation and the global graph representation, but differ in their way of generating negative instances.
@@
Second, to sample the positive or negative instances, different augmentation techniques have been proposed for learning robust representations. The most straightforward augmentation strategy perturbs or samples the graphs in a uniformly random fashion Veličković et al. (2018); Zhu et al. (2020), while recent studies turn to a more strategic selection of augmentations. On one hand, MVGRL Hassani and Khasahmadi (2020) applies graph diffusion to transform the adjacency matrix as structure-space augmentation. On the other hand, JOAO You et al. (2021) presents an automated framework that exploits the most challenging augmentation relative to the current loss.
Method Reconstruction objective
node feature node degree edge adj. matrix graph feature
VGAE Kipf and Welling (2016) $\times$ $\times$ $\times$ $\checkmark$ $\times$
GPT-GNN Hu et al. (2020b) $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
LaGraph Xie et al. (2022) $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
GraphMAE Hou et al. (2022) $\checkmark$ $\times$ $\times$ $\times$ $\times$
GraphMAE2 Hou et al. (2023) $\checkmark$ $\times$ $\times$ $\times$ $\times$
Table 3: Summary of generative pre-training strategies.

### v4

@@ IV-A Standard Meta-learning Techniques
Formally, in the meta-training phase, consider a set of meta-training tasks $\mathcal{T}_{\text{train}}=\{\mathcal{T}^{1}_{\text{train}},\mathcal{T}^{2}_{\text{train}},\ldots,\mathcal{T}^{n}_{\text{train}}\}$. Each task $\mathcal{T}^{i}_{\text{train}}=(\mathcal{S}^{i}_{\text{train}},\mathcal{Q}^{i}_{\text{train}})\in\mathcal{T}_{\text{train}}$ consists of a support set $\mathcal{S}^{i}_{\text{train}}$ and a query set $\mathcal{Q}^{i}_{\text{train}}$. Subsequently, the process of meta-learning, also called learning-to-learn [31], involves optimizing the loss over the set of meta-training tasks $\mathcal{T}^{i}_{\text{train}}$, as follows.
$\omega^{*}=\arg\min_{\omega}\mathbb{E}_{\mathcal{T}^{i}_{\text{train}}\in\mathcal{T}_{\text{train}}}\mathcal{L}(\mathcal{T}^{i}_{\text{train}};\omega),$ (4)
where $\omega^{*}$ represents the prior or meta-knowledge extracted by the meta-learner from the meta-training tasks, describing “how to learn”. The meta-knowledge $\omega^{*}$ can then be transferred to downstream novel tasks, i.e., training a model for the downstream task under the guidance of $\omega^{*}$. Specifically, in the meta-testing phase, consider a set of meta-testing tasks $\mathcal{T}_{\text{test}}=\{\mathcal{T}^{1}_{\text{test}},\mathcal{T}^{2}_{\text{test}},\ldots,\mathcal{T}^{m}_{\text{test}}\}$, where each task $\mathcal{T}^{i}_{\text{test}}=(\mathcal{S}^{i}_{\text{test}},\mathcal{Q}^{i}_{\text{test}})\in\mathcal{T}_{\text{test}}$ also includes a few-shot support set $\mathcal{S}^{i}_{\text{test}}$ and an unlabeled query set $\mathcal{Q}^{i}_{\text{test}}$. Given the extracted meta-knowledge $\omega^{*}$ and a meta-testing task $\mathcal{T}^{i}_{\text{test}}=(\mathcal{S}^{i}_{\text{test}},\mathcal{Q}^{i}_{\text{test}})\in\mathcal{T}_{\text{test}}$, meta-learning trains a specialized downstream model for task $\mathcal{T}^{i}_{\text{test}}$ by optimizing the loss on its support set $\mathcal{S}^{i}_{\text{test}}$:
$\theta^{*}_{i}=\arg\min_{\theta}\mathcal{L}(\mathcal{S}^{i}_{\text{test}},\omega^{*};\theta),$ (5)
where $\theta$ denotes the parameters of the downstream model. The optimized parameters $\theta^{*}_{i}$ can then be used for prediction on the query set $\mathcal{Q}^{i}_{\text{test}}$ of the task $\mathcal{T}^{i}_{\text{test}}$.
Toward few-shot learning on graphs, these standard meta-learning methods can be enhanced in various ways. We categorize the literature based on their graph-specific enhancements within the meta-learning framework into two types: structure-based enhancements and adaptation-based enhancements, as summarized in Table IV and Table V, respectively.
IV-B Structure-based Enhancement on Graphs
@@
Node-level enhancement. The nodes of a graph encompass implicit structural information for enhancing the meta-learning model. Recent studies predominantly focus on differentiating node weights in a task to reflect their varying structural importance. Specifically, in Protonets-based methods such as GPN [27] and FAAN [28], it is observed that each support node contributes variably to the prototype of each class in a task. Thus, they propose to evaluate the importance of nodes in the support set and adjust their contribution w.r.t. the query node when constructing the prototypes.
@@
Edge-level enhancement. Edges are fundamental to graphs as they represent explicit first-order structures. Moreover, a sequence of edges forms a path, which is capable of representing more complex high-order interactions among distant nodes. As a result, many meta-learning methods on graphs incorporate edge- or path-based enhancements. On one hand, edge-based enhancement leverages auxiliary information associated with edges, thereby extending beyond mere node-level information. For instance, to boost few-shot link prediction in knowledge graphs, HMNet [34] adopts matching networks (MN) as its meta-learning framework, performing matching for both entities and relations. In particular, for relation-level matching, it calculates a relation-aware score using a relation encoder to enrich the link prediction task. Another work [143] develops a generative model to generate new triplets, serving to augment the sparse links within a MAML-based framework. On the other hand, paths can capture long-range dependencies between distant nodes. RALE [89] proposes to exploit the paths between each query node and the support nodes for task-level dependencies, and between each target node and the hub nodes for graph-level dependencies. Furthermore, MetaHIN [25] employs meta-paths [144] to explore heterogeneous semantic relationships between users and items, aiming to acquire finer-grained semantic priors for new tasks.
@@
Subgraph-level enhancement. Subgraphs provide rich contextual information to its containing nodes. Hence, recent studies have utilized subgraphs to enhance the expressiveness of nodes. G-Meta [66] generates class prototypes based on the contextual subgraph of each support node, and similarly expands each query node to its subgraph, to enhance the robustness of prototype and query representations. Similarly, GEN [93] proposes a graph extrapolation network to extrapolate knowledge through the neighbors of the support set, which essentially exploits the one-hop subgraphs of the support nodes. Moreover, meta-tail2vec [132] introduces the notion of locality-aware tasks, where the support set is sampled from the neighborhood subgraph of the query node, leveraging the structural proximity between support and query nodes.
TABLE IV: Structure-based meta-learning enhancement for few-shot learning on graphs.
Method Structure Meta-learner Task
enhancement Node Edge Graph
GPN [27] node Protonets $\checkmark$ $\times$ $\times$
FAAN [28] node Protonets $\times$ $\checkmark$ $\times$
HMNet [34] edge MN $\times$ $\checkmark$ $\times$
[143] edge MAML $\times$ $\checkmark$ $\times$
RALE [89] path MAML $\checkmark$ $\times$ $\times$
MetaHIN [25] path MAML $\checkmark$ $\times$ $\times$
G-Meta [66] subgraph hybrid $\checkmark$ $\checkmark$ $\times$
GEN [93] subgraph Protonets $\times$ $\checkmark$ $\times$
meta-tail2vec [132] subgraph MAML $\checkmark$ $\checkmark$ $\times$
IV-C Adaptation-based Enhancement on Graphs
@@
Graph-wise adaptation. GFL [36] recognizes the topological variances across different graphs, and thus customizes a global prior for each individual graph. More concretely, it generates class prototypes specifically tailored to each graph, utilizing a gate function based on graph-specific representation that is applied to the global prior. MI-GNN [145] adopts a similar idea and employs a Feature-wise Linear Modulation (FiLM) [146] to modulate the global prior for each graph before undergoing a further task-wise adaptation within a MAML framework.
Task-wise adaptation. Meanwhile, there is potential for enhancement in task-wise adaptation. In the context of multi-label few-shot classification, MetaTNE [147] observes an increase in task diversity due to the multi-label setting. Thus, adopting a uniform node embedding across such diverse tasks may result in inadequate expressiveness, given that the same node could be associated with different labels in different tasks. To address this, MetaTNE conducts an additional adaptation for the node embeddings w.r.t. the query set in each task, thereby better adapting to the characteristics of the task. Moreover, variation in feature distribution across tasks can significantly impair the performance of meta-learning. To mitigate this issue, AMM-GNN [65] capitalizes on the FiLM mechanism [146] to modulate the feature matrix for each task. It customized a task-specific feature matrix, enabling a globally shared GNN model to apply these tailored features for task-wise adaptation in few-shot node classification.
Others. In addition to enhancing graph- or task-wise adaptation, several approaches enhances downstream adaptation in various other ways. AS-MAML [148] improves adaptation from an optimization standpoint. It employs a reinforcement learning-based controller to determine the optimal step size for the adaptation process in MAML. MetaDyGNN [137] introduces hybrid adaptations for dynamic graphs that entails both time- and node-wise adaptations within a MAML framework to solve few-shot link prediction problems.
TABLE V: Adaptation-based meta-learning enhancement for few-shot learning on graphs.
Method Adaptation Meta Task
enhancement learner Node Edge Graph
GFL [36] graph MAML $\checkmark$ $\times$ $\times$
MI-GNN [145] graph hybrid $\checkmark$ $\times$ $\times$
MetaTNE [32] task Protonets $\checkmark$ $\times$ $\times$
AMM-GNN [65] task MAML $\checkmark$ $\times$ $\times$
AS-MAML [148] step size MAML $\times$ $\times$ $\checkmark$
MetaDyGNN [137] hybrid MAML $\times$ $\checkmark$ $\times$
IV-D Discussion
@@
In summary, existing research for few-shot learning on graphs often enhances a standard meta-learner, either through structural augmentation or by refining the adaptation process. However, these approaches share a common drawback inherent to meta-learning approaches. First, they still require abundant labels for a base set during the meta-training phase, which can be challenging or costly to acquire. Second, they fail to leverage the vast amount of unlabeled data to learn a more comprehensive prior. Third, they are limited by the i.i.d. assumption in task distribution, and cannot handle different types of tasks downstream. This leads to a further challenge: Can we address a diverse range of few-shot tasks on graphs without an extensively annotated base set, while utilizing abundant unlabeled graphs? Recently, pre-training on graphs has emerged as a promising solution to this challenge.
V Pre-training Techniques on Graphs
@@
Given the abundance of unlabeled data in many domains and the flexibility to deal with many types of downstream tasks, pre-training approaches have been popularized [37]. In graph context, the pre-training stage typically utilizes self-supervised methods on unlabeled graph data to train a graph encoder, which aims to capture task-agnostic intrinsic properties of graphs, such as node features and local or global structures. Subsequently, the pre-trained graph encoder, serving as a form of prior knowledge, can be used to address a variety of downstream tasks through an adaptation stage. An overview of the pre-training and adaptation process is depicted in Fig. 6.
In this section, we provide a taxonomy of existing work in this direction, covering both pre-training and adaptation.
@@
TABLE VI: Summary of contrastive graph pre-training.
Method Instance Augmentation Graph types
GRACE [72] node uniform general
GCC [30] graph uniform general
GraphCL [40] graph uniform general
SimGRACE [74] graph perturbing encoder general
GraphLoG [73] dataset uniform general
DGI [29] cross-scale uniform general
InfoGraph [42] cross-scale uniform general
Subg-Con [71] cross-scale uniform general
MVGRL [149] cross-scale diffusion general
JOAO [41] graph adaptive to loss general
GCGM [150] node adaptive to loss general
You et al. [151] graph view generator general
GCA [152] node adaptive to instance general
HeCo [153] node uniform hetero.
CPT-HG [154] cross-scale uniform hetero.
PT-HGNN [155] cross-scale uniform hetero.
SelfRGNN [76] node curvature over time dynamic
DDGCL [156] graph uniform dynamic
CPDG [75] cross-scale temporal-aware sampling dynamic
GearNet [157] graph uniform 3D
V-A Strategies for the Pre-training Stage
@@
Current self-supervised tasks (also known as pretext tasks) for graph pre-training predominantly fall into two broad categories: contrastive and generative strategies, although many approaches employ both strategies simultaneously to extract more comprehensive knowledge from graphs. More recently, some graph pre-training approaches have begun to leverage the power of pre-trained large language models.
V-A1 Contrastive Strategies
@@
Contrastive pre-training functions by contrasting instances at various scales within a graph. For each target instance, contrastive methods sample positive and negative instances, aiming to bring the positive instances closer to the target in the latent space while distancing the negative instances, as shown in Fig. 6(b). Formally, for each target instance $o$ in the pre-training data $\mathcal{T}_{\text{pre}}$, the contrastive task samples or constructs a set of positive instances $\mathcal{P}_{o}$, along with a set of negative instances $\mathcal{N}_{o}$. The contrastive loss function can be generally formulated as
$\textstyle-\sum_{o\in\mathcal{T}_{\text{pre}}}\ln\frac{\sum_{a\in\mathcal{P}_{o}}\exp\left(\frac{\mathtt{sim}(\mathbf{h}_{a},\mathbf{h}_{o})}{\tau}\right)}{\sum_{a\in\mathcal{P}_{o}}\exp\left(\frac{\mathtt{sim}(\mathbf{h}_{a},\mathbf{h}_{o})}{\tau}\right)+\sum_{b\in\mathcal{N}_{o}}\exp\left(\frac{\mathtt{sim}(\mathbf{h}_{b},\mathbf{h}_{o})}{\tau}\right)}$ (6)
where $\mathtt{sim}(\cdot,\cdot)$ represents a similarity function such as cosine.
We summarize contrastive strategies for graph pre-training in Table VI, and discuss them based on their instance scale, augmentation techniques, and types of graph.
@@
Instance scale. First, the scale of instances used for contrasting varies, ranging from individual nodes to broader subgraphs, or even the entire graph. One one hand, node-level approaches provide a more granular treatment focusing on the representations of each individual node. An example is GRACE [72], which generates two augmented views by randomly corrupting the original graph. Then, for each target node in one view, it treats the corresponding node from the other view as positive, while others as negative. On the other hand, graph-level approaches expand the scope to (sub)graphs, thereby capturing global structural patterns. GCC [30] conducts random walk in an $r$-hop ego network to obtain a target subgraph instance alongside its positive instances. Negative instances are acquired by performing random walk in an $r^{\prime}$-hop ego network of the same ego node, for some $r\neq r^{\prime}$. Similarly, GraphCL [40] generates graph views with various data augmentations and treats graph views augmented from the same graph as positive instances, while those from different graphs as negative instances. Extending from the graph level to the dataset level, GraphLoG [73] introduces hierarchical prototypes to capture not only the local-instance structure among correlated graphs, but also the global-semantic structure across all graphs in the dataset. Moreover, there also exist cross-scale approaches contrasting between node- and graph-level instances, such as DGI [29] and InfoGraph [42]. They both maximize the mutual information between the local node representation and the global graph representation, but differ how they generate negative instances. On the other hand, Subg-Con [71] captures local structure information by contrasting central nodes with their contextual subgraphs for improved scalability on very large graphs.
@@
Augmentation. To sample the positive or negative instances, different augmentation techniques have been proposed for learning robust representations. The most straightforward augmentation strategy perturbs or samples the graphs in a uniformly random fashion [29, 72], while recent studies turn to a more strategic selection of augmentations. MVGRL [149] applies graph diffusion to transform the adjacency matrix as structure-space augmentation. Moreover, JOAO [41] and GCGM [150] attempt to automatically select the most challenging augmentations adapted to the current loss, while GCA [152] adapts each graph’s augmentation according to the importance of node or edge features. On the other hand, et al. [151] propose a learnable generator for graph augmentation. In contrast, SimGRACE [74] does not apply any augmentation to the original graph; instead, it perturbs the graph encoder to produce two correlated views.
Graph types. In addition to general homogeneous graphs, contrastive methods are also widely applied across a diverse range of graph types, including heterogeneous graphs, dynamic graphs and 3D graphs [158, 159]. For heterogeneous graphs, HeCo [153] contrasts between a schema view and a meta-path view to capture complex semantics through local and high-order structures on a heterogeneous graph. CPT-HG [154] propose two contrastive strategies at the relation level and subgraph level. The former considers two types of negative samples from inconsistent relations and disconnected nodes, respectively, while the latter leverage meta-graphs [160] to construct positive and negative samples. PT-HGNN [155] proposes a dual contrastive mechanism focusing on node and schema level tasks. Node level negative samples are pairs of nodes that match specific relations, while schema level negative samples are generated under the guidance of the network schema. For dynamic graphs, DDGCL [156] treats a time-weighted subgraph view of a target node as postive samples, whereas taking its structurally shuffled temporal subgraph and unweighted subgraph views as negative. Furthermore, SelfRGNN [76] proposes a Riemannian GNN with time-varying curvature, and further leverages the functional curvature over time to obtain auxiliary views without generating new graphs as augmentations. Additionally, to learn long-short term evolution patterns and distinct structural patterns in dynamic graphs, CPDG [75] designs a structural-temporal sampler with temporal-aware sampling probability to extract more meaningful contextual subgraphs as positive and negative instances. For 3D graph structures, GearNet [157] pre-trains protein representations using their 3D structures, positing that biologically relevant substructures are positioned near each other, while those that are unrelated are mapped farther apart.
TABLE VII: Summary of generative graph pre-training.
Method Reconstruction objective Graph type
node feat. node deg. edge adj. matrix graph feat. other info.
VGAE [43] $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ general
GPT-GNN [39] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ general
MaskGAE [77] $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ general
NWR-GAE [161] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ general
LaGraph [162] $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ general
GraphMAE [163] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ general
GraphMAE2 [78] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ general
Liu et al. [164] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ KG
Wen et al. [79] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ KG
MPKG [165] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ KG
PT-DGNN [166] $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ dynamic
STEP [167] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ dynamic
PMGT [168] $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ MMG
ColdGPT [169] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ MMG
V-A2 Generative Strategies